OpenStudy (anonymous):

A line passes through a point (-2,4), and its y intercept is (0,6). What is the equation of the line that is perpendicular to the first line and passes through the point (5,-4)

OpenStudy (anonymous):

\(\LARGE\color{pink}{Welcome~to~Open~Study!!!}\)

OpenStudy (anonymous):

First, find the slope of the first line m = (6-4)/(0--2) = 2/2 = 1 We want a line that is perpendicular to the first so we will want a slope that is the opposite reciprocal, so the slope of our second line will be (-1). Now we can use point-slope form to write the equation of the second line containing the point (5, -4) and a slope of (-1) y - (y1) = m(x - (x1)) y - (-4) = -1(x - (5)) y + 4 = -x + 5 y = -x +1